OpenStudy (anonymous):

how do you solve (x-y)(x^2-11xy+y^2)

OpenStudy (mathstudent55):

What do you mean by "solve"? Solving is usually done to equations. There is no equal sign here, so you don't have an equation.

OpenStudy (anonymous):

it tells me to find the product how do you find it?

OpenStudy (mathstudent55):

Ok. You are multiplying a binomial by a trinomial. Multiply every term of the binomial by every term of the trinomial. Then combine like terms.

OpenStudy (mathstudent55):

For example: \((2x + 3)(3x^2 + 4x - 5)\) \(=2x \times 3x^2 + 2x \times 4x + 2x \times (-5) + 3 \times 3x^2 + 3 \times 4x + 3 \times (-5)\) \(= 6x^3 + 8x^2 - 10x + 9x^2 + 12x - 15\) \(= 6x^3 + 17x^2 + 2x - 15\)

OpenStudy (anonymous):

i did that but it tell me it was not right

OpenStudy (mathstudent55):

Can you show your steps? I'll see if you did anything wrong.